# Reversible Decoupling Network for Single Image Reflection Removal

### v1

@@ 3.2 Transmission-rate-Aware Prompt Generator
Previous methods for SIRR often exhibit limited generalization capabilities due to the inherent complexity and variability of optical factors in real-world reflective scenarios, compounded by the constraint of limited training data. This limitation can be observed in the real-world test samples we collected, as shown in Fig. 4. Meanwhile, in both real-world and synthesized data, color/intensity is often compromised due to the reflection overlaying the transmission, with the transmission $T$ itself being degraded by a transmission rate $a$. In image restoration tasks, the ground truths are typically clean images. But, linearly deviated input/result often occurs because of color/illumination shifts in the real-world scenarios. The phenomenon is further detailed in the appendix A.
To solve the aforementioned problems, we develop a transmission-rate estimator using a simplified version of the ConvNext model (Liu et al., 2022) pre-trained on ImageNet-1k (Deng et al., 2009). Given an input image $I\in\mathbb{R}^{3\times H\times W}$, our transmission-rate estimator predicts six parameters: $\alpha_{\{R,G,B\}},\beta_{\{R,G,B\}}$ such that $\|\alpha_{i}T+\beta_{i}-I\|_{2}$ is minimized for each $i\in\{R,G,B\}$. When testing the input image using the six parameters generated by the prompt generator, we can obtain an average PSNR of 24.34dB across four benchmark datasets (Real20, Objects, Postcard, and Wild), surpassing the previous state-of-the-art method by Dong et al. (Dong et al., 2021). This result confirms the effectiveness of our estimated transmission rate.
Once the transmission rate factor $\alpha_{\{R,G,B\}},\beta_{\{R,G,B\}}$ is estimated, a three-layer MLP is used to generate prompts that guide the MRCE, resulting in a prompt $P\in\mathbb{R}^{C\times H\times W}$, where $C$ represents the output dimension of the patch embedding layer, set to 64 in our work. Subsequently, the prompt is used to modulate the intermediate features from the column embedding layer $F$ into $P\circ F$, which allows the network to dynamically adapt to the specific characteristics of each input image, thereby enhancing the accuracy of reflection removal.
@@ 3.3 Training Objective
Our model undergoes two training stages. In the first stage, we train the estimator for the transmission rate. Once this is complete, we fix the classifier and proceed to train the main model along with the prompt generator. This training scheme ensures that both the transmission-rate-aware prompt generator and the main model work harmoniously towards the task, resulting in a robust solution.
@@
Content Loss. The content loss ensures consistency between the output images and the ground truth training data. In the image domain, we adopt the Mean Squared Error (MSE) loss. Following previous works (Hu & Guo, 2023; 2021), we further regularize the model by encouraging consistency between the output and ground truth in the gradient domain, which writes:
@@
Perceptual Loss. To enhance the perceptual quality of images produced by our model, we minimize, we minimize the $\ell_{1}$ discrepancy between the features of predicted elements and the ground-truth references. This comparison is made at the ‘conv2_2’, ‘conv3_2’, ‘conv4_2’, and ‘conv5_2’ layers of a pre-trained VGG-19 network on the ImageNet dataset. Denoting the features at the $i$th layer as $\phi_{i}(\cdot)$, the perceptual loss is computed as:
@@
where $\omega_{j}$ are weighting coefficients for each layer. The total loss turns out to be:
@@ 4.1 Implementation Details
Our model is implemented in PyTorch (Paszke et al., 2019) and optimized with Adam optimizer (Kingma & Ba, 2015) on an RTX 3090 GPU for 20 epochs. The learning rate is initialized at $10^{-4}$, and remains fixed throughout the training phase, with a batch size of 2. The training dataset comprises both real and synthetic images. To align with previous works, we evaluate the performance of our model under two commonly used data settings: a) The setting from (Hu & Guo, 2021; Wei et al., 2019) and (Li et al., 2023), which consists of 90 real image pairs from (Zhang et al., 2018) and 7,643 synthesized pairs from the PASCAL VOC dataset (Everingham et al., 2010); and b) The setting from (Hu & Guo, 2023) and (Dong et al., 2021), which includes 200 additional real image pairs provided by (Li et al., 2020). For data synthesizing, we follow the pipeline and physical model from DSRNet (Hu & Guo, 2023), represented by $I=\alpha T+\beta R-T\circ R$. Slightly, we modify this approach by sampling individual $\alpha$ and $\beta$ for R, G, and B channels. This adjustment aims to prevent the transmission rate estimator from converging to a trivial solution. The parameters the PHE are initialized by a pretrained FocalNet (Yang et al., 2022).
@@
Table 2: Quantitative results on the Nature dataset. The competitors are all trained with the additional data from the Nature dataset
ERRNet IBCLN YTMT DSRNet Zhu et al. Ours
PSNR 22.18 23.57 23.85 25.22 26.04 26.21
SSIM 0.756 0.783 0.810 0.832 0.846 0.842
@@ 4.2 Performance Evaluation
For the comparison, we evaluate seven state-of-the-art methods: ERRNet (Wei et al., 2019), IBCLN (Li et al., 2020), RAGNet (Li et al., 2023), Dong et al. (Dong et al., 2021), YTMT (Hu & Guo, 2021), DSRNet (Hu & Guo, 2023), Zhu et al. (Zhu et al., 2024), on four real-world datasets, including Real20 (Zhang et al., 2018) and three subsets of the $\text{SIR}^{2}$ Datasets (Wan et al., 2017), for the Nature Dong et al. (2021) dataset, we compare IBCLN, ERRNet, YTMT, DSRNet and Zhu et al..
Quantitative comparisons. The quantitative result is shown in Tab. 1. We directly employ the code and pre-trained weights publicly provided by their authors to obtain all the quantitative results. To make a fair comparison, the methods with and without additional data from the Nature dataset are compared separately. Apparently, our methods show their superiority over other competitors on all testing datasets, only falling short on SSIM compared to Zhu et al. on the Objects dataset. Our methods achieved a promising boost, especially on the Real20 dataset, which contains hard cases collected in real-world conditions, meaning our method can better fit real-world conditions. The other three datasets contain a variety of scenes, illumination conditions, and glass thickness, meaning our method performs better in most conditions. The experimental result demonstrates that our proposed SIRS method can adapt to complicated situations and has a stronger generalization ability.
For a comprehensive comparison, we present the results obtained on the Nature dataset in Tab. 2, which comprises 20 real-world samples. Our method achieved the best PSNR and the second-best SSIM, with a marginal decrease of only 0.004 in SSIM. These results further underscore the superiority of our approach in real-world scenarios.
@@
Figure 3: Qualitative comparisons on samples from the Wild dataset. Please zoom in for more details. More visual results can be found in the appendix.
(a) Input
(b) ERRNet
(c) IBCLN
(d) Dong et al.
(e) YTMT
(f) DSRNet
(g) Zhu et al.
(h) Ours
Qualitative comparisons. The qualitative comparison is shown in Fig. 3 and Fig. 4, with additional visual examples provided in the appendix. The first case in Fig. 3 illustrates a highly reflective object, which presents a significant challenge for reflection removal techniques due to its intensity. Our method successfully eliminates the reflective object, accurately revealing the underlying texture and color information. This performance is superior to other methods, highlighting our approach’s effectiveness in handling complex real-world reflections. In contrast, ERRNet, RAGNet, Dong et al., YTMT, DSRNet and Zhu et al. struggle to remove the object, leaving it almost entirely intact. Although IBCLN partially removes the reflection, it fails to recover the underlying color information, resulting in an incomplete outcome. This example clearly demonstrates our method’s advanced capability in accurately identifying and removing even strong and complex reflections, further proving its robustness in real-world scenarios.
The second example further showcases our method’s proficiency in handling reflections spread across an image. Here, the reflection is complex and covers a large area, which other methods fail to remove effectively. In contrast, our approach accurately targets and eliminates the majority of the reflection, preserving the integrity of the non-reflective elements.
Figure 4 illustrates the robustness of our method in real-world scenarios. These two cases were captured in real-life conditions by us. In the first example, a dense reflection covers the car window, a challenge that competing methods largely fail to address, with only Zhu et al. managing partial removal. However, our approach almost entirely separates the reflections, producing more visually appealing results. A similar outcome is observed in the second example, where our method successfully removes nearly all reflections. In contrast, all other methods struggle to handle this scenario effectively. These examples demonstrate the robustness of our decoupling paradigm, confirming its effectiveness in real-world scenarios.
These results demonstrate the effectiveness of our decoupling routine, offering several key advantages: 1) accurate identification and separation of reflection components from underlying content, 2) robust performance in removing dense reflections common in real-world scenarios, and 3) strong generalizability across diverse conditions. Collectively, these findings validate the theoretical soundness and practical efficacy of our proposed method.
@@ 4.3 Ablation Studies
To better verify the effect of our prompt generator and reversible network structure, we bring a series of ablation studies, including different settings of network structure and prompt generator. The results are gathered in Tab. 3. We present the results of our prompt generator on the left side and the results of our network structure on the right side.
Table 3: Ablation studies on the prompt generator and different network configurations.
Setting Prompt Pre. Average Setting Dual-stream Ref. Loss Invertibility Average
PSNR SSIM PSNR SSIM
A $\times$ $\times$ 25.52 0.909 D $\checkmark$ $\checkmark$ $\checkmark$ 26.37 0.917
B $\times$ $\checkmark$ 25.99 0.910 E $\times$ $\times$ $\checkmark$ 25.99 0.914
C $\checkmark$ $\checkmark$ 26.03 0.913 F $\times$ $\checkmark$ $\times$ 24.05 0.884
Ours $\checkmark$ $\times$ 26.65 0.917 Ours $\times$ $\checkmark$ $\checkmark$ 26.65 0.917
@@
Discussion on transmission-rate-aware prompt generator. To inform the model with the transmission rate, a straightforward approach is to adjust the input image to enhance it globally using the estimated transmission rate. Specifically, for $I:=aT+bR+\phi(T,R)$, we adjust the input $I$ to $\frac{1}{a}I:=T+\frac{b}{a}R+\frac{1}{a}\phi(T,R)$. This operation is denoted as Pre. in Tab. 3. As shown in Tab. 3, if we remove all transmission-rate-aware techniques (setting A), the average performance drops by 1.13 dB. If we adopt the straightforward method described above (setting B), the performance recovers by 0.47 dB. This confirms the importance of informing the model with the transmission rate.
However, as we analyzed in Section 3.2, directly adjusting the input image is far from optimal. Due to potential inaccuracies in the estimation in some scenarios, directly adjusting the model can introduce an additional shift that is difficult to correct during second-stage training. A more subtle and flexible approach is to reweight the feature channels with our transmission-rate-aware prompt.
To verify this, we both adjust the input and add a transmission-rate-aware prompt to the feature (setting C). The performance remains nearly the same as in Setting B, indicating that adjusting the input makes it challenging for the model to recover from incorrect estimations. Finally, our model with the proposed transmission-rate-aware prompt outperforms all variants, demonstrating its efficacy.
@@
Discussion on model design. To verify the rationality of our design of the decoupling model, we created three new variants of our model. We modify our RDNet to a DSRNet-style one, where two streams estimate transmission and reflection separately in a single column, and interact with each other. This variant is denoted as Dual-stream (Setting D). As shown in Tab. 3, even with double computation, the performance still drops by 0.28 dB. This confirms the superiority of our decoupling design compared to the dual-stream design. Secondly, we removed the reflection part ($c_{1}\|\hat{R}-R\|^{2}_{2}$) in the content loss function (Eq. (5)), leaving only the transmission part ($c_{0}\|\hat{T}-T\|^{2}_{2}+c_{2}\|\nabla\hat{T}-\nabla T\|_{1}$) in the training process. This variant is denoted as Ref. Loss. (Setting E). A performance drop of 0.66dB can be observed. This confirms the necessity of the reflection loss function. Without regularization predicting the other component, the network weakens its ability to clearly identify both components in single-stream feature maps.
To verify the necessity of the invertibility of the network in the reflection removal task, we replace the reversible connection with the U-Net connection (Ronneberger et al., 2015) (Setting F). Although it requires slightly more parameters and much more memory, a massive performance drop of 2.6 dB can be discovered, indicating the importance of invertibility design.
@@ 5 Conclusion
In this paper, we proposed RDNet, a novel model for addressing key challenges in the task of single image reflection removal. Specifically, RDNet tackles the limitations of insufficient utilization of multi-scale, pretrained hierarchical information and information loss during feature decoupling. The multi-column reversible structure enables the preservation of rich semantic features, which are then effectively leveraged in the multi-scale processing of each column. Furthermore, the proposed Transmission-rate-Aware Prompt Generator alleviates the inherent conflict between complex reflection parameters and limited training data. Through these innovations, RDNet demonstrates an enhanced capability for robust reflection removal. Our method demonstrates superior performance compared to state-of-the-art techniques across a range of real-world benchmark datasets, highlighting its robustness and adaptability in diverse reflective scenarios. Ablation studies further validate the effectiveness of our key contributions, confirming the advantages of our design choices. It is positive that our work opens up new avenues for research in reflection removal, and has the potential to impact various applications in computer vision and image processing significantly.
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant nos. 62372251 and 62072327. Mingjia Li is also supported by Cloud TPUs from TPU Research Cloud (TRC).
Appendix
Appendix A Further discussion on the TAPG
Figure 5: Visualization of the drawback of Mean Squared Error (MSE). Both color shifts and noise degradation exhibit the same MSE relative to the ground truth.
As illustrated in Fig. 5 with a toy visualization of a pure white image, even though the color bias and noise exhibit the same Mean Squared Error (MSE) relative to the ground truth, a linear estimation can instantly correct the image’s color shift, whereas noise requires more complex operations to address. Metrics like MSE and Mean Absolute Error (MAE) struggle to compel the network to effectively recognize and rectify the linear degradation in the physical formulations. In this context, by pre-calibrating the features with a transmission-rate-aware prompt, we can significantly mitigate the effects of linear degradation, such as color and intensity inconsistencies.
Appendix B Qualitative Comparisons
More visual cases. We exhibit a total of nine additional cases: two cases from the Real20 dataset in Fig. 9, three cases from the Solid dataset in Fig.7, two cases from the Postcard dataset in Fig. 8 and two real-world cases captured by us in Fig. 6. As illustrated, our method excels at revealing the information obscured by reflections and is highly effective in removing the majority of the reflections.
Table 4: The experiment of changing numbers of columns. The best results are indicated in bold.
Num Col Average
PSNR SSIM
2 26.25 0.914
4 26.65 0.917
6 26.19 0.910
Appendix C Additional Experiments
The ablation study for the number of columns. In this study, we investigate the effect of varying the number of columns on the overall performance in Tab. 4. Specifically, we adjusted the number of columns after the first PHE column, experimenting with configurations of 2, 4, and 6 columns. Our findings indicate that a configuration with 4 columns yields the highest performance. In contrast, configurations with 2 and 6 columns resulted in performance drops of 0.4dB and 0.46dB in PSNR, respectively. This suggests that an optimal balance exists, where too few or too many columns can detract from the model’s performance.
(a) Input
(b) ERRNet
(c) IBCLN
(d) Dong et al.
(e) YTMT
(f) DSRNet
(g) Zhu et al.
(h) Ours
Figure 6: Visual comparison of estimated transmission layers between state-of-the-arts and ours on real-world samples.
(a) Input
(b) ERRNet
(c) IBCLN
(d) RAGNet
(e) Dong et al.
(f) YTMT
(g) DSRNet
(h) Zhu et al.
(i) Ours
(j) GT
Figure 7: Visual comparison of estimated transmission layers between state-of-the-arts and ours on Objects dataset.
(a) Input
(b) ERRNet
(c) IBCLN
(d) RAGNet
(e) Dong et al.
(f) YTMT
(g) DSRNet
(h) Zhu et al.
(i) Ours
(j) GT
Figure 8: Visual comparison of estimated transmission layers between state-of-the-arts and ours on Postcard dataset.
(a) Input
(b) ERRNet
(c) IBCLN
(d) RAGNet
(e) Dong et al.
(f) YTMT
(g) DSRNet
(h) Zhu et al.
(i) Ours
(j) GT
Figure 9: Visual comparison of estimated transmission layers between state-of-the-arts and ours on real-world samples (Real 20).

### v2

@@ 3.2 Transmission-rate-Aware Prompt Generator
Previous methods for SIRR often exhibit limited generalization capabilities due to the inherent complexity and variability of optical factors in real-world reflective scenarios, compounded by the constraint of limited training data. This limitation can be observed in the real-world test samples we collected, as shown in Fig. 3. Meanwhile, in both real-world and synthesized data, color/intensity is often compromised due to the reflection overlaying the transmission, with the transmission $T$ itself being degraded by a transmission rate $a$. In image restoration tasks, the ground truths are typically clean images. However, linearly deviated input/result occurs due to color/illumination shifts in real-world scenarios.
To solve the aforementioned problems, we develop a transmission-rate estimator using a simplified version of the ConvNext model [27] pre-trained on ImageNet-1k [5]. Given an input image $I\in\mathbb{R}^{3\times H\times W}$, our transmission-rate estimator predicts six parameters: $\alpha_{\{R,G,B\}},\beta_{\{R,G,B\}}$ such that $\|\alpha_{i}T+\beta_{i}-I\|_{2}$ is minimized for each $i\in\{R,G,B\}$. When testing the input image using the six parameters generated by the prompt generator, we can obtain an average PSNR of 24.34dB across four benchmark datasets (Real20, Objects, Postcard, and Wild), surpassing the previous state-of-the-art method by Dong et al. [9]. This result confirms the effectiveness of our estimated transmission rate.
Once the transmission rate factor $\alpha_{\{R,G,B\}},\beta_{\{R,G,B\}}$ is estimated, a three-layer MLP is used to generate prompts for MRCE, resulting in a prompt $P\in\mathbb{R}^{C\times H\times W}$, where $C$ represents the output dimension of the patch embedding layer, set to 64 in our work. Subsequently, the prompt is used to modulate the intermediate features from the column embedding layer $F$ into $P\circ F$, which allows the network to dynamically adapt to the specific characteristics of each input image, thus enhancing the accuracy of reflection removal.
@@ 3.3 Training Objective
Our model undergoes two training stages. In the first stage, we train the estimator for the transmission rate. Once this is complete, we fix the classifier and proceed to train the main model along with the prompt generator. This training scheme ensures that both the transmission-rate-aware prompt generator and the main model work harmoniously toward the task, resulting in a robust solution.
@@
Content Loss. The content loss ensures consistency between the output images and the ground truth training data. In the image domain, we adopt the Mean Squared Error (MSE) loss. Following previous works [15, 14], we further regularize the model by encouraging consistency between the output and ground truth in the gradient domain, which writes:
@@
Perceptual Loss. To enhance the perceptual quality of images produced by our model, we minimize the $\ell_{1}$ discrepancy between the features of predicted elements and the ground-truth references. This comparison is made at the ‘conv2_2’, ‘conv3_2’, ‘conv4_2’, and ‘conv5_2’ layers of a pre-trained VGG-19 network on the ImageNet dataset. Denoting the features at the $i$th layer as $\phi_{i}(\cdot)$, the perceptual loss is computed as:
@@
where $\omega_{j}$ are weighting coefficients for each layer and $\omega_{j}$ $=0.2$ is empirically set. The total loss turns out to be:
@@ 4.1 Implementation Details
Our model is implemented in PyTorch [29] and optimized with Adam optimizer [19] on an RTX 3090 GPU for 20 epochs. The learning rate is initialized at $10^{-4}$, and remains fixed throughout the training phase, with a batch size of 2. The training dataset comprises both real and synthetic images. To align with previous works, we evaluate the performance of our model under two commonly used data settings: a) The setting from [14, 43] and [24], which consists of 90 real image pairs from [49] and 7,643 synthesized pairs from the PASCAL VOC dataset [10]; and b) The setting from [15] and [9], which includes 200 additional real image pairs provided by [21]. For data synthesizing, we follow the pipeline and physical model from DSRNet [15], represented by $I=\alpha T+\beta R-T\circ R$. Slightly, we modify this approach by sampling individual $\alpha$ and $\beta$ for R, G, and B channels. This adjustment aims to prevent the transmission rate estimator from converging to a trivial solution. The parameters of the PHE are initialized by a pretrained FocalNet [47].
@@ 4.2 Performance Evaluation
For the comparison, we evaluate seven state-of-the-art methods: ERRNet [43], IBCLN [21], RAGNet [24], Dong et al. [9], YTMT [14], DSRNet [15], Zhu et al. [54], on four real-world datasets, including Real20 [49] and three subsets of the $\text{SIR}^{2}$ Datasets [39], for the Nature [9] dataset, we compare IBCLN, ERRNet, YTMT, DSRNet and Zhu et al..
Setting Prompt Adjusting Input PSNR SSIM
A $\times$ $\times$ 25.52 0.909
B $\times$ $\checkmark$ 25.99 0.910
C $\checkmark$ $\checkmark$ 26.03 0.913
Ours $\checkmark$ $\times$ 26.65 0.917
Table 2: Ablation studies on the prompt generator. The PSNR and SSIM are calculate across Real20, Objects, Postcard and Wild.
Setting Dual-stream Ref. Loss Invertibility PSNR SSIM
D $\checkmark$ $\checkmark$ $\checkmark$ 26.37 0.917
E $\times$ $\times$ $\checkmark$ 25.99 0.914
F $\times$ $\checkmark$ $\times$ 24.05 0.884
Ours $\times$ $\checkmark$ $\checkmark$ 26.65 0.917
Table 3: Ablation studies for network configurations. The PSNR/SSIM are calculated on Real20, Objects, Postcard, and Wild.
Competitor DSRNet Zhu et al.
Win-rate of our method 64.4% 77.2%
Table 4: User study comparison with DSRNet and Zhu et al.
Quantitative comparisons. The quantitative result is shown in Tab. 1. We directly employ the code and pre-trained weights publicly provided by their authors to obtain all the quantitative results. To make a fair comparison, the methods with and without additional data from the Nature dataset are compared separately. Apparently, our methods show their superiority over other competitors on all testing datasets, only falling short on SSIM compared to Zhu et al. on the Objects dataset. Our methods achieved a promising boost, especially on the Real20 dataset, which contains hard cases collected in real-world conditions, meaning our method can better fit real-world conditions. The other three datasets contain a variety of scenes, illumination conditions, and glass thickness, meaning our method performs better in most conditions. The experimental result demonstrates that our method can adapt to complicated situations and has a stronger generalization ability. For a comprehensive comparison, we present the results obtained on the Nature dataset in supplementary, which comprises 20 real-world samples. Our method achieved the best PSNR and the second-best SSIM, with a marginal decrease of only 0.004 in SSIM. These results further underscore the superiority of our approach in real-world scenarios.
@@
Figure 4: Qualitative comparisons on samples from the Wild dataset. Please zoom in for more details.
Qualitative comparisons. The qualitative comparison is shown in Fig. 4 and Fig. 3, with additional examples in the supplementary. Fig. 4 presents a challenging case with a highly reflective object. Our method effectively removes the reflection, revealing clear texture and color information, outperforming other approaches. In comparison, other methods struggle to remove the object fully. This example highlights our method’s robustness in complex real-world scenarios. The second example further showcases our method’s proficiency in handling reflections spread across an image. Here, the reflection is complex and covers a large area, which other methods fail to remove effectively. In contrast, our approach accurately targets and eliminates the majority of the reflection, preserving the integrity of the non-reflective elements. Figure 3 demonstrates the robustness of our method in real-world scenarios, with examples captured under natural conditions. In the first case, a dense reflection covers the car window, posing a significant challenge that most competing methods fail to overcome—only Zhu et al. achieves partial removal. In contrast, our approach effectively separates the reflections, producing clearer and more visually appealing results. Similarly, in the second example, our method successfully removes nearly all reflections, while other methods struggle to perform effectively. These cases underscore the robustness of our decoupling approach.
User Study. We present our user study in Tab. 4. Inspired by LIME-Eval[22], we asked 20 users to choose their favorable one from each pair of outputs. All five datasets are included and 587 pairs of options are obtained. Our method achieves a win rate of 77.2% to Zhu et al., and 64.4% to DSRNet, indicating better performance.
@@ 4.3 Ablation Studies
To better verify the effect of our prompt generator and reversible structure, we provide a series of ablation studies for the key component of our design as below.
@@
Discussion on transmission-rate-aware prompt generator. To inform the model with the transmission rate, a straightforward approach is to adjust the input image to enhance it globally using the estimated transmission rate. Specifically, for $I:=aT+bR+\phi(T,R)$, we adjust the input $I$ to $\frac{1}{a}I:=T+\frac{b}{a}R+\frac{1}{a}\phi(T,R)$. This operation is denoted as Adjusting Input in Tab. 2. As shown in Tab. 2, if we remove all transmission-rate-aware techniques (setting A), the average performance drops by 1.13 dB. If we adopt the straightforward method described above (setting B), the performance recovers by 0.47 dB. This confirms the importance of informing the model with the transmission rate. However, as we analyzed in Section 3.2, directly adjusting the input image is far from optimal. Due to potential inaccuracies in the estimation in some scenarios, directly adjusting the model can introduce an additional shift that is difficult to correct during second-stage training. A more subtle and flexible approach is to reweight the feature channels with our transmission-rate-aware prompt. To verify this, we both adjust the input and add a transmission-rate-aware prompt to the feature (setting C). The performance remains nearly the same as in Setting B, indicating that adjusting the input makes it challenging for the model to recover from incorrect estimations. Finally, our model with the proposed transmission-rate-aware prompt outperforms all variants, demonstrating its efficacy.
@@
Discussion on model design. To verify the rationality of our design of the decoupling model, we created three new variants of our model. We modify our RDNet to a DSRNet-style one, where two streams estimate transmission and reflection separately in a single column and interact with each other (Setting D). As shown in Tab. 3, even with double computation, the performance still drops by 0.28 dB. This confirms the effectiveness of the decoupling design compared to the dual-stream design. Secondly, we removed the reflection part ($c_{1}\|\hat{R}-R\|^{2}_{2}$) in the content loss function (Eq. 8), leaving only the transmission part ($c_{0}\|\hat{T}-T\|^{2}_{2}+c_{2}\|\nabla\hat{T}-\nabla T\|_{1}$) in the training process. This variant is denoted as Setting E. A performance drop of 0.66dB can be observed. This confirms the necessity of the reflection loss function. Without regularization predicting the other component, the network weakens its ability to clearly identify both components in single-stream feature maps. To verify the necessity of the invertibility of the network, we replace the reversible connection with the U-Net connection [31] (Setting F). With slightly more parameters and much more memory, a massive performance drop of 2.6 dB can be discovered, indicating the importance of invertibility.
Column count 2 4 6
PSNR 26.25 26.65 26.19
SSIM 0.914 0.917 0.910
Table 5: The experiment of changing numbers of columns.
Discussion on the number of columns. We investigate the effect of the number of columns in Tab.5. Specifically, we adjusted the number of columns with configurations of 2, 4, and 6 columns. Our findings indicate that our choice of 4 columns yields the highest performance. In contrast, configurations with 2 and 6 columns resulted in performance drops of 0.4dB and 0.46dB in PSNR, respectively. This suggests that an optimal balance exists, where too few or too many columns can detract from the model’s performance.
@@ 5 Conclusion
In this paper, we proposed RDNet, a novel model for addressing key challenges in the task of single-image reflection removal. Specifically, RDNet tackles the limitations of insufficient utilization of multi-scale, pretrained hierarchical information and information loss during feature decoupling. The multi-column reversible structure enables the preservation of rich semantic features, which are then effectively leveraged in the multi-scale processing of each column. Furthermore, the proposed Transmission-rate-Aware Prompt Generator alleviates the inherent conflict between complex reflection parameters and limited training data. Through these innovations, RDNet demonstrates an enhanced capability for robust reflection removal. Our method demonstrates superior performance compared to state-of-the-art techniques across real-world benchmark datasets, highlighting its robustness and adaptability in diverse reflective scenarios. It is positive that our work opens up new avenues for research in reflection removal and has the potential to significantly impact various applications in image restoration.
Acknowledgement. This work was supported by the National Natural Science Foundation of China under Grant nos.62372251 and 62072327. The computational resources of this work is partially supported by TPU Research Cloud (TRC) and DataCanvas Limited.